# Newtonian potential from scattering amplitudes in super-renormalizable gravity

### v1

@@ Abstract
Conclusion. We have investigated the Newtonian potential in super-renormalizable gravity through scattering amplitudes, systematically including both tree-level and one-loop contributions in $D$ dimensions. At tree level, the potential in $D=4$ is nonsingular and governed by an error function, thereby regularizing the classical divergence. We further performed the first one-loop calculation of the $\mathcal{O}(G^{2})$ correction, which required deriving new Feynman rules for the three-graviton vertex in the nonlocal framework. Although the one-loop result involves intricate integrals that preclude simple analytic forms, a numerical evaluation in four dimensions confirms unambiguously that the potential remains finite at the origin. This provides the first conclusive evidence that singularity resolution in super-renormalizable gravity holds at the one-loop level. Whether this remarkable property persists at higher loop orders remains an important question for future investigation.
@@ Acknowledgements.
This study was supported by the National Natural Science Foundation of China (Grant No. 12333008).
@@ Appendix B Triangle Loop Integrals
We now turn to the triangle integrals governing the one-loop correction, shown in Fig. 2, beginning with their form in GR. Following the definitions in Eqs. (LABEL:28)-(LABEL:32), we first analyze the fundamental structure of these integrals by temporarily setting aside explicit vertex factors. The simplest and most illustrative case is the scalar triangle integral
@@
where the integral discussed above belong to the category of nonlocal integrals, making the conventional Wick rotation in QFT inapplicable. However, an alternative method can be employed to handle such integral Buoninfante:2022krn . Specifically, the integration variable $l_{0}$ can be defined on the imaginary axis, which is equivalent to condition $l_{0}=il_{D}$ ($l^{2}=-l^{2}_{E}$). Subsequently, the scattering amplitude calculated on the imaginary axis is mapped to the physical conclusion on the real axis.
@@
Now, we focus on the integrals associated with the contributions from nonlocal terms, which are categorized into three types as defined in Eq. (LABEL:29). Using the same approach, we directly present the final results for the integrals involved.

### v2

@@ IV The Metric correction from One-Loop Scattering Amplitudes
V Summary and conclusion
In this work, we investigate the modifications to the Newtonian potential in $D$-dimensional spacetime within the framework of super-renormalizable gravity, focusing on the first- and second-order corrections to the gravitational constant $G_{N}$ through the Feynman scattering amplitudes approach. We first introduce the super-renormalizable gravity, analyzing it from two key perspectives: (1) from the viewpoint of the propagator and unitarity, where we observe that unitarity is preserved at the tree-level. and (2) from the power-counting perspective, where we identify that the primary source of divergence arises from the one-loop diagrams, thereby further confirming the super-renormalizable.
Subsequently, we examine the Newtonian potential through the Feynman scattering amplitudes, focusing on both tree-level and loop-level contributions. At tree level, the potential in $D=4$ is nonsingular and governed by an error function, thereby regularizing the classical divergence. We further performed the first one-loop calculation of the $\mathcal{O}(G^{2})$ correction, which required deriving new Feynman rules for the three-graviton vertex in the nonlocal framework. Although the one-loop result involves intricate integrals that preclude simple analytic forms, a numerical evaluation in four dimensions confirms unambiguously that the potential remains finite at the origin. This provides the first conclusive evidence that singularity resolution in super-renormalizable gravity holds at the one-loop level. Whether this remarkable property persists at higher loop orders remains an important question for future investigation.
@@ Acknowledgements.
This study was supported by the National Natural Science Foundation of China (Grant No. 12333008) and the Basic Research Program of Shanxi Province (Grant No. 202503021211204).
@@ Appendix B Triangle Loop Integrals
We now turn to the triangle integrals governing the one-loop correction, shown in Fig. 2, beginning with their form in GR. Following the definitions in Eqs. (28)-(LABEL:33), we first analyze the fundamental structure of these integrals by temporarily setting aside explicit vertex factors. The simplest and most illustrative case is the scalar triangle integral
@@
where the integral discussed above belong to the category of nonlocal integrals, making the conventional Wick rotation in QFT inapplicable. However, an alternative method can be employed to handle such integral Buoninfante (2022). Specifically, the integration variable $l_{0}$ can be defined on the imaginary axis, which is equivalent to condition $l_{0}=il_{D}$ ($l^{2}=-l^{2}_{E}$). Subsequently, the scattering amplitude calculated on the imaginary axis is mapped to the physical conclusion on the real axis.
@@
Now, we focus on the integrals associated with the contributions from nonlocal terms, which are categorized into three types as defined in Eq. (LABEL:33). Using the same approach, we directly present the final results for the integrals involved.